Please help me with this word problem.
The base of a ladder is 14 feet away from the wall. The top of the ladder is 17 feet from the floor. Find the length of the ladder to the nearest thousandth.

Thank you!

Answer by checkley71(8403)

HERE WE HAVE A RIGHT TRIANGLE WITH SIDES OF 14 & 17 FEET & NEED TO FIND THE HYPOTENUSE WHICH IS C IN THE PYTHAGOREAN THERUM.
C^2=A^2+B^2
C^2=14^2+17^2
C^2=196+289
C^2=485
C=SQRT485
C=22.02 IS THE LENGTH OF THE HYPOTENUSE.